The following data were collected during a short race between two friends. Velocity (m/s) 0 0.5 1 1.5 2 2 4 6 2 0 Time (s) 0 2 4 6 8 10 12 14 16 18 a) Describe the different sections of the graph. b) Determine the acceleration over the first eight seconds. c) Determine the maximum acceleration. d) Using the graph calculate the displacement: i) over the first eight seconds ii) the total race. e) Find the maximum velocity reached by the runner.



1
Expert's answer
2021-12-06T16:42:36-0500
"a=0.25\frac{m}{s^2}\\a_m=1\frac{m}{s^2}\\s=2\ m\\S=0\ m\\v_m=6\frac{m}{s}"